# Open Problems, and Future Trends

### v1

@@ V-B Privacy and Security Issues
In Section III, we have extensively discussed the confidentiality and security issues of the Industrial Metaverse from the perspectives of data, network, distributed platforms, and enabling technologies, as well as existing solutions. It can be foreseen that, as the Metaverse platform gradually matures, its security measures will largely rely on continuous upgrades through security patches. However, with the increasing prevalence of sophisticated cyber attacks, such security measures may bring challenges in terms of operational costs and scalability[149]. In the future, it is possible to adopt intrinsic security solutions, such as quantum key distribution, which utilizes channel-based keys generated through quantum entanglement to address information leakage in wireless transmissions[102]. Relevant skills/mathematical tools: Quantum key distribution, cryptography, network security.
@@ V-C Lightweight and Ease of Use
Since Industrial Metaverse scenarios are constantly producing, storing, and transmitting data, they require massive access to hardware devices. These devices include production equipment deployed in factories, industrial sensor devices, and access devices such as VR glasses for user interaction. The lightweighting at the hardware level has been slow to materialize. Ways to reduce hardware overhead as well as increase the flexibility and scalability of devices is a valuable topic[150]. Relevant skills/mathematical tools: Embedded system design, hardware optimization, sensor technology.
@@ V-D Model Accuracy and Completeness
The key factor affecting model accuracy in assembly line mode manufacturing applications is the quality of the data collected. The user must provide the system with complete and accurate available data[151]. Although the addition of DT technology can realize virtual samples directly into the model, the data sources often have serious heterogeneity due to different equipment models and data types, and the data fusion and calibration methods at this stage still need to be improved. Relevant skills/mathematical tools: Industrial data preprocessing methods, data fusion methods, machine learning algorithms.
@@ V-E Unified Architecture
Central to the realization of the Metaverse lies the crucial foundation of Web 3.0 architecture. However, the intricate task of amalgamating diverse technologies within a unified architecture necessitates comprehensive efforts. In order to address this challenge, the development of common protocols capable of accommodating the distinct needs and requirements of various industrial sectors becomes imperative. Such protocols would serve as the bedrock for seamless integration, enabling interoperability and synergy among different technologies. By fostering a standardized framework, the development of the Metaverse can progress cohesively, accommodating a multitude of industries and catering to the unique demands of users across diverse domains[152]. Relevant skills/mathematical tools: Network protocol design, distributed system architecture, Web technologies.
V-F Components are Fully Decentralized
The core feature of Web3.0 is decentralization, but APIs at this stage do not integrate well with Web3.0, and new decentralized common protocols need to be developed to achieve complete decentralization of Metaverse components[153]. Only the successful implementation of a decentralized network will occur. This is a challenge because web components and servers still run through legacy controls and protocols, and therefore the backbone infrastructure will need to be extensively changed[154]. Therefore, in the scope of the future, the design of decentralized protocols running on the web needs to be investigated, and open standards and APIs need to be integrated with Web 3.0 engines where data can be seamlessly transferred to different endpoints. Relevant skills/mathematical tools: Decentralized technologies[155], BC development, smart contract programming.
V-G Efficient Cross-chain Authentication
@@
Efficient cross-chain authentication and governance is crucial to ensure the security and legitimacy of digital asset-related activities (e.g., asset transactions) across different submetadomains built on heterogeneous BCs. Current cross-chain mechanisms are mainly focused on digital asset transfers and rely on notary schemes, hash locks, relay chains and side chains, with few mechanisms considering cross-chain authentication and governance for meta-domains. The implementation, efficiency and security of cross-domain and BC authentication need further research. Moreover, novel decentralized, hierarchical and penetrating cross-chain governance mechanisms also need further research in the Metaverse. In addition, efficient meta-space-specific consensus mechanisms, redesigned block structures, and well-designed user incentives are all necessary for unique meta-space applications[156]. In summary, open challenges include application-specific governance rule design, programmable and scalable cross-chain governance architecture design, on-chain entity identification and risk assessment, dynamic and collaborative cross-chain governance, etc. Relevant skills/mathematical tools: BC cross-chain technology, authentication protocols, smart contract programming.
V-H Integration with Large-scale Pre-trained Models
@@
Large-scale pre-trained models hold the potential to offer exquisitely tailored and individualized encounters within the Metaverse. Through discerning analysis of user inclinations, behaviors, and historical data, these models can be refined to generate bespoke virtual panoramas[157], characters, and interactions, ushering in a realm of heightened immersion and personalized engagement within the Metaverse.
@@
Nevertheless, notable challenges persist. One such challenge lies within the realm of the language macrocosm, where Metaverse users hail from diverse nations and regions, conversing in various languages. Consequently, it becomes imperative for the model to possess the capability to seamlessly interact across linguistic boundaries. Furthermore, the endeavor to ensure the seamless deployment of the model, endowed with high concurrency and unwavering stability, becomes an ineluctable quandary[158]. This ensures swift and responsive handling of users’ requests, a matter of utmost significance. Relevant skills/mathematical tools: Natural language processing, deep learning model training, model deployment and optimization.
V-I Integration with Quantum Computing
@@
Quantum computing, with its unparalleled capacity to handle intricate calculations and tackle problems exponentially faster, possesses the potential to reshape the capabilities of the Metaverse and elevate its performance to unprecedented heights. A recent study conducted by Cui et al.[159] delves into the implementation of secure communication and efficient cross-chain protocols utilizing quantum computing. Meanwhile, Ren et al.[160] have explored the utilization of game-based quantum collective learning and many-to-many matching schemes within the Metaverse, enabling the attainment of optimal strategies for maximizing system revenue. Looking ahead, quantum computing is poised to find widespread application in enhancing data processing, encryption, and security performance within the Metaverse. Furthermore, it holds the promise of expediting virtual and AR experiences, thereby propelling the Metaverse into a realm of heightened immersion and seamless interactivity. Relevant skills/mathematical tools: Quantum computing theory, quantum algorithm design, quantum communication protocols.
V-J Availability of Datasets for Industrial Metaverse
@@
In industrial environments, there is a plethora of data in different types, formats, and sources. These data often reside in disparate systems and devices, leading to data isolation and fragmentation. Lack of unified data integration and management mechanisms hampers seamless integration and consolidation of data. Moreover, various stakeholders in industrial environments may possess different data ownership and access rights. Establishing appropriate mechanisms for data ownership and access control is essential to ensure data legitimacy and compliance. Relevant skills/mathematical tools: Data management, data integration, access control.
V-K Ethical and Social Impacts
When machine decisions impact people’s lives, especially in industrial scenarios with high cost and safety requirements, it is important to clarify how much autonomy the system allocates to machine algorithms. In addition to the technical aspects, fairness in resource allocation must also be considered[161]. For example, some large companies may find it easier to concentrate high-quality and extensive digital resources for DT modeling, while expensive hardware may exclude organizations with lower purchasing power, exacerbating the digital divide. It is necessary to enact more artificial intelligence laws in the future to subject the operation of the industrial Metaverse to legal jurisdiction.

### v2

@@ V-B Privacy and Security Issues
In Section III, we have extensively discussed the confidentiality and security issues of the Industrial Metaverse from the perspectives of data, network, distributed platforms, and enabling technologies, as well as existing solutions. It can be foreseen that, as the Metaverse platform gradually matures, its security measures will largely rely on continuous upgrades through security patches. However, with the increasing prevalence of sophisticated cyber attacks, such security measures may bring challenges in terms of operational costs and scalability[190]. In the future, it is possible to adopt intrinsic security solutions based on knowledge in the fields of quantum key distribution, cryptography, and network security, such as quantum key distribution, which utilizes channel-based keys generated through quantum entanglement to address information leakage in wireless transmissions.
@@ V-C Lightweight and Ease of Use
Since Industrial Metaverse scenarios are constantly producing, storing, and transmitting data, they require massive access to hardware devices. These devices include production equipment deployed in factories, industrial sensor devices, and access devices such as VR glasses for user interaction. The lightweighting at the hardware level has been slow to materialize. Ways to reduce hardware overhead as well as increase the flexibility and scalability of devices is a valuable topic[191]. Therefore, how to achieve effective support for these requirements through refined embedded system design, optimized hardware performance, and improved sensor technology is the direction that researchers need to focus on.
@@ V-D Model Accuracy and Completeness
The key factor affecting model accuracy in assembly line mode manufacturing applications is the quality of the data collected. The user must provide the system with complete and accurate available data[192]. Although the addition of DT technology can realize virtual samples directly into the model, the data sources often have serious heterogeneity due to different equipment models and data types, and the data fusion and calibration methods at this stage still need to be improved. To better improve the performance of the models, researchers need to acquire skills related to industrial data preprocessing, data fusion and machine learning algorithms.
@@ V-E Unified Architecture
Central to the realization of the Metaverse lies the crucial foundation of Web 3.0 architecture. However, the intricate task of amalgamating diverse technologies within a unified architecture necessitates comprehensive efforts in web protocol design, distributed systems architecture, and web technologies. In order to address this challenge, the development of common protocols capable of accommodating the distinct needs and requirements of various industrial sectors becomes imperative. Such protocols would serve as the bedrock for seamless integration, enabling interoperability and synergy among different technologies. By fostering a standardized framework, the development of the Metaverse can progress cohesively, accommodating a multitude of industries and catering to the unique demands of users across diverse domains[193].
V-F Decentralization in Industrial Systems
Although systems in the Industrial Metaverse are typically not fully decentralized—some distributed components are often owned and managed by central authorities—in certain industrial scenarios, such as edge computing and multi-party collaboration platforms, decentralization can be used to reduce latency or enhance system resilience. The core feature of Web3.0 is decentralization, but current APIs are not well integrated with Web3.0. Therefore, in specific scenarios (such as supply chain management, distributed energy management systems, etc.), new decentralized protocols need to be developed to achieve the decentralization of Metaverse components[194]. While some components can be deployed in a distributed manner, the existing infrastructure relies on traditional controls and protocols, which means that to achieve decentralization, significant changes to the backbone architecture are necessary[195]. Therefore, the focus of future research is on designing decentralized protocols suitable for the Web and integrating open standards and APIs with Web 3.0 engines to enable seamless data transfer to different endpoints[196].
V-G Network Isolation Issue
Although OT (Operational Technology) and IT (Information Technology) networks have been integrated, they still exist in isolated silos. OT and IT systems typically use different technology stacks and protocols. OT systems are mostly specialized industrial control systems, while IT systems focus more on information storage, processing, and exchange. These technological differences create significant challenges in interoperability, data transmission, and management when integrating OT and IT networks. Although Szucs et al.[197] provide a use case that collects data from OT/IT systems and prevents interface confusion, thus simplifying the enterprise network architecture, the issues of coordinating between different technology stacks, protocols, and security requirements, as well as the specific implementation details, remain to be addressed.
V-H Efficient Cross-chain Authentication
@@
Efficient cross-chain authentication and governance is crucial to ensure the security and legitimacy of digital asset-related activities (e.g., asset transactions) across different submetadomains built on heterogeneous BCs. Current cross-chain mechanisms are mainly focused on digital asset transfers and rely on notary schemes, hash locks, relay chains and side chains, with few mechanisms considering cross-chain authentication and governance for meta-domains. The implementation, efficiency and security of cross-domain and BC authentication need further research. Moreover, novel decentralized, hierarchical and penetrating cross-chain governance mechanisms also need further research in the Metaverse. In addition, efficient meta-space-specific consensus mechanisms, redesigned block structures, and well-designed user incentives are all necessary for unique meta-space applications[198]. Achieving these goals requires a deep understanding of BC cross-chain technologies, authentication protocols, and the ability to program smart contracts in order to design and implement scalable cross-chain governance architectures, dynamically collaborative governance rules, and on-chain entity identification and risk assessment mechanisms.
V-I Integration with Large-scale Pre-trained Models
@@
Large-scale pre-trained models hold the potential to offer exquisitely tailored and individualized encounters within the Metaverse. Through discerning analysis of user inclinations, behaviors, and historical data, these models can be refined to generate bespoke virtual panoramas, characters, and interactions, ushering in a realm of heightened immersion and personalized engagement within the Metaverse.
@@
Nevertheless, notable challenges persist. One such challenge lies within the realm of the language macrocosm, where Metaverse users hail from diverse nations and regions, conversing in various languages. Consequently, it becomes imperative for the model to possess the capability to seamlessly interact across linguistic boundaries. Additionally, ensuring the seamless deployment, high concurrency, and stability of the model is essential to quickly respond to and handle user requests[199]. Addressing these challenges requires expertise in natural language processing, deep learning model training, and skills in model deployment and optimization to support multilingual interactions and efficient user engagement within the Metaverse.
V-J Integration with Quantum Computing
@@
Quantum computing, with its unparalleled capacity to handle intricate calculations and tackle problems exponentially faster, possesses the potential to reshape the capabilities of the Metaverse and elevate its performance to unprecedented heights. A recent study conducted by Cui et al.[200] delves into the implementation of secure communication and efficient cross-chain protocols utilizing quantum computing. Meanwhile, Ren et al.[201] have explored the utilization of game-based quantum collective learning and many-to-many matching schemes within the Metaverse, enabling the attainment of optimal strategies for maximizing system revenue.Looking ahead, quantum computing is poised to find widespread application in enhancing data processing, encryption, and security performance within the Metaverse. Furthermore, it holds the promise of expediting virtual and AR experiences, thereby propelling the Metaverse into a realm of heightened immersion and seamless interactivity. Advancing these innovations requires a strong grasp of quantum computing theory, quantum algorithm design, and quantum communication protocols.
V-K Availability of Datasets for Industrial Metaverse
@@
In industrial environments, there is a large amount of data of different types, formats, and sources, which are often distributed across various systems and devices, leading to data isolation and fragmentation. The lack of unified data integration and management mechanisms hinders the seamless consolidation and utilization of data. Moreover, different stakeholders in industrial settings may have varying data ownership and access rights. Therefore, establishing appropriate data ownership and access control mechanisms based on theories of data management, data integration, and access control is crucial to ensuring data legitimacy and compliance.
V-L Ethical and Social Impacts
When machine decisions impact people’s lives, particularly in industrial environments with high demands on cost and safety, it is crucial to define the level of autonomy granted to algorithms within the system. Beyond technical considerations, the ethical use of data and the fair distribution of rights in the Industrial Metaverse must also be addressed, especially in terms of corporate social responsibility [202]. Large corporations are more likely to concentrate digital resources for advanced digital twin modeling, while the high cost of hardware may exclude smaller enterprises with limited budgets, thus exacerbating the digital divide.
With the advancement of automation technologies, AI-driven robots and virtual simulation systems will significantly reduce the need for human operators and traditional on-site workers. While this enhances production efficiency, it could also lead to the gradual obsolescence of physical roles such as machine operators, technicians, and on-site inspectors. Moreover, as industrial processes shift to digital environments, workers in traditional industries may face challenges in adapting to new skill demands, such as operating VR-based simulations, AI analytics, and digital twin technologies.
In the B2C context, protecting user privacy, data security, and combating misinformation are of critical importance. Companies should also promote diversity and inclusion, focus on environmental sustainability, and strike a balance between innovation and social responsibility[203]. Further AI regulations should be introduced to ensure that the Industrial Metaverse operates within legal frameworks, safeguarding users’ rights.